# SpikingBrain: Spiking Brain-inspired Large Models

### v2

@@ 2.2 Integrated Model Architectures
The 7B model achieves purely linear complexity by interleaving linear attention and sliding window attention (SWA) layers with a fixed 4K window in a 1:1 ratio. The FFN modules adopt the same SwiGLU design as the base model. Spike coding is applied to convert the activations of all linear projection layers in the model, as matrix multiplications for parameter projections dominate the model’s FLOPs. In this architecture, the inter-layer hybrid design is optimized for extreme efficiency, owing to its simplicity in attention computation, model parallelism and sequence parallelism. SWA layers captures precise local patterns, while linear attention layers efficiently compresses long-range information. The model thus provides linear-time training complexity and constant memory usage during inference, regardless of sequence length, offering substantial efficiency gains for long-context processing. In practice, we implement linear attention using a Gated Linear Attention Module yang2024gated (GLA), where the gating vector $\mathbf{g}_{t}$, derived from low-rank projection parameters, enhances expressivity and recurrent modeling capability:
@@
The 76B model integrates linear attention and SWA in a 1:1 intra-layer parallel hybrid configuration, while standard full-attention layers are interleaved at a 1:6 ratio across layers. This intra-layer hybrid design is optimized for stronger language modeling capabilities, as it enables more intricate attention computation and incorporates additional attention parameters and FLOPs. During parallel hybridization, outputs of both attention branches are normalized to ensure consistent scale and to prevent instability in the early stages of training. In addition, 128 learnable sink tokens (xiao2023efficient, ; dong2024hymba, ) are introduced to mitigate the attention-sink phenomenon in softmax attention and to enhance the flexibility of SWA for local modeling. Specifically, 128 trainable embeddings are prepended to the input; these tokens are attended by all other tokens and also attend to each other without causal masking. This functionality is implemented by customizing FlashAttention (daoflashattention, ) kernels. The model also employs Gated Linear Attention Modules, but in this case, the gating vector is tied directly to the key vector (chou2024metala, ; qinhgrn2, ), i.e.,
$\mathbf{g}_{t}=\mathbf{1}-\mathbf{k}_{t},$ (12)
eliminating the need for extra gating parameters. The FFN modules adopt a sparse Mixture-of-Experts (MoE) design: each MoE layer consists of 16 routed experts (top-1 activated) and one shared expert, so that only about 15% of parameters (12B) are activated per token. To stabilize training and control parameter growth, 7 dense FFN layers are preserved out of 28 total layers (specifically at layers 1, 2, 3, 5, 7, 9, and 11). Similarly, spike coding is applied to the activations of all linear projection layers throughout the model.
@@
Our architectural choices are closely aligned with principles observed in biological brains. i) Linear attention modules exhibit modeling properties analogous to human memory, relying on compressed and continuously updated states (gu2025tradeoffs, ). At each step, they retrieve information only from the current memory, showing Markov-like behavior. From a biological perspective, their stateful temporal recurrence can be viewed as a simplified abstraction of dendritic dynamics with multi-branch morphology. ii) The Mixture-of-Experts (MoE) component embodies the principle of modular sparse activation and functional specialization, reminiscent of the distributed and specialized processing found in neural circuits (o2021and, ). iii) Our spike coding scheme draws inspiration from event-driven and adaptively sparse neuronal activation in biological systems (maass1997networks, ). By combining network-level sparsity (MoE) with neuron-level spiking sparsity, our approach enables on-demand allocation of computation and provides a robust two-scale efficiency strategy. Collectively, these results suggest a promising pathway for designing large-model architectures that are both efficient and biologically plausible (10636118, ).
@@ 3.1 Attention Map Correspondence
On this basis, SWA can be interpreted as a sparsified version of this map with a strong recency bias (beltagy2020longformer, ; child2019generating, ; zaheer2020big, ):
@@
Moreover, linear attention can be regarded as a low-rank approximation of the attention map, where the maximum rank of $\mathbf{A}$ is $d$ (the dimensionality of the query/key vectors) (chen2021scatterbrain, ; wang2020linformer, ):
@@
Leveraging this attention-map correspondence, we can initialize the QKV projection parameters of SpikingBrain directly from a pre-trained Transformer checkpoint (kasai2021finetuning, ), and this applies to full attention, linear attention, and SWA. By reusing the learned QK similarity and performing lightweight training on a small amount of data, the attention can be adapted to these local or low-rank cases. Since SWA maintains local modeling precision and linear attention retains global interaction, combining them in a hybrid paradigm provides a closer approximation to the original attention map (dass2023vitality, ; chen2021scatterbrain, ; arora2024simple, ). This enables a smoother transition during conversion and accelerates loss convergence.
@@
  • Apply non-negative activation to the QK vectors in linear attention (kasai2021finetuning, ; mercat2024linearizing, ), such as ReLU or Sigmoid. Because softmax attention maps are inherently non-negative, the converted attention must preserve this property to enable smooth transfer. Normalization, as in softmax attention, is applied after the linear attention outputs.
@@
  • Fully train the model during conversion to ensure performance, using learning rate warm-up and either cross-architecture distillation (wang2024mamba, ; zhang2024lolcats, ) or full-parameter training (zhang2024gated, ). For simplicity and to fit the memory capacity of MetaX GPUs, we adopt full-parameter training without freezing the backbone.
@@ 3.2 MoE Upcycling
To efficiently expand a dense model into an MoE architecture, we adopt the upcycling technique (komatsuzaki2022sparse, ), which increases model capacity while reusing the knowledge encoded in the original parameters. At initialization, the FFN in the base dense model is replicated into $N$ experts, and a randomized router is introduced. The router selects the top-$k$ experts for each token with probability $\mathbf{p}$ and outputs their weighted sum, ensuring functional equivalence to the original dense FFN at the start:
@@
Directly replicating and activating multiple experts, however, amplifies the output scale. To maintain consistency between MoE and dense FFN outputs, we rescale the expert weights (he2024upcycling, ). The relationship between MoE activation and dense activation is:
$\mathrm{MoE\_activation}$ $=S\times\mathrm{dense\_activation}+\frac{k}{N}\mathrm{dense\_activation}$ (19)
$=(S+\frac{k}{N})\mathrm{dense\_activation}.$ (20)
Accordingly, the scaling factor is defined as:
$\text{scaling\_factor}=\sqrt[3]{\frac{1}{S+\frac{k}{N}}}.$ (21)
We apply this factor to both shared and routed experts at initialization:
$E_{i}(\mathbf{x}):=\text{scaling\_factor}\times\mathrm{FFN}(\mathbf{x}),\quad\text{at initialization}.$ (22)
@@ 3.3.1 Step 1: Adaptive Threshold Spiking
$\mathbf{v}_{t+1}$ $=\mathbf{v}_{t}-V_{\text{th}}\cdot\mathbf{s}_{t}+\mathbf{x}_{t+1},$ (24)
$\mathbf{s}_{t}$ $=1\ \text{ if }\ \mathbf{v}_{t}\geq V_{\text{th}}(\mathbf{x})\ \text{ else }\ 0.$ (25)
@@
This single-step integer formulation enables computationally efficient and stable optimization on GPUs, while the adaptive threshold maintains appropriate firing rates. As a result, our spiking scheme preserves the essential characteristics of biological neurons, yet is sufficiently simplified to be practical for large-scale models.
For gradient computation in the optimization phase, since the spiking process is essentially an approximate representation of the original membrane potential, the gradient of the spiking process can be skipped (i.e., using straight-through estimation), which maintains the gradient computation process consistent with conventional training.
@@
  • Threshold dynamics: $V_{th}(\mathbf{x})$ is determined by the mean absolute membrane potential. When the input potential is high, $V_{th}(\mathbf{x})$ increases, preventing excessive spiking and thus controlling sparsity to reduce redundant computation. When the input potential is low, $V_{th}(\mathbf{x})$ decreases, allowing neurons to emit a small number of spikes to retain key information and avoid accuracy loss from inactivity. Statistically, membrane potentials often follow a long-tailed Gaussian-like distribution with occasional outliers. In this setting, the mean absolute value approximates 0.8 times the standard deviation, providing a stable metric for regulating spike activity.
  • Effect of the hyperparameter $k$: $k$ directly scales the threshold and thereby determines the distribution of spike counts. A larger $k$ lowers $V_{th}(\mathbf{x})$, producing higher spike counts $\mathbf{s}_{INT}$ and broader firing ranges. This is suited for accuracy-critical scenarios where additional computation is acceptable. A smaller $k$ raises $V_{th}(\mathbf{x})$, reducing spike counts and producing sparser encodings, which is advantageous for low-power edge deployments. Tuning $k$ thus provides a flexible trade-off between accuracy and efficiency.
@@ 3.3.2 Step 2: Spike Coding
During inference, the integer spike count $\mathbf{s}_{\text{INT}}$ generated in Step 1 is expanded into a sparse spike sequence with values $\{0,1,-1\}$ along the time dimension to support event-driven computation. The process is formulated as:
@@
  • Binary Spike Coding {0,1}: This is the most basic event-driven spike coding method. Each time a unit spike is fired (with a value of $1$), it represents an activation of the neuron state, and the spike count is accumulated over continuous time steps. This coding scheme is intuitive and simple, with low computational overhead, making it suitable for scenarios with very low spike counts and effectively reducing system complexity. However, when representing large counts, this coding often requires many time steps to complete the representation. Additionally, the lack of directional optimization for spike firing results in a higher firing rate, further limiting system energy efficiency. Meanwhile, this encoding scheme only supports positive integers (unsigned $\mathbf{s}_{\text{INT}}$) which is not applicable to Adaptive Threshold Spiking.
  • Ternary Spike Coding {-1,0,1}: To improve neuronal expressivity and sparsity in event-driven computations, we introduce ternary spike coding. The core idea is to extend the traditional binary spike coding by adding inhibitory spikes (1-1), resulting in three firing states: {-1,0,1}. Compared to binary coding, which can only represent positive values, ternary coding offers bidirectional expression capabilities. Here, "11" represents excitatory spikes, "1-1" inhibitory spikes, and "0" the silent state, making it more aligned with the "excitation/inhibition" regulatory mechanism in biological neural systems. As shown in Figure 4 (b), ternary spike coding not only provides symmetric expression capabilities of $\pm 1$ but also reconstructs the mapping from membrane potential to spike counts through a symmetric quantization strategy. This strategy maps activations to a symmetric distribution $k,\dots,0,\dots,+k]$ rather than $0,1,2,\dots]$, so that low-amplitude counts (e.g., $\pm 1$) statistically occupy a larger probability mass, effectively absorbing high-frequency large counts from the tail of the original distribution. As a result, this scheme halves the number of time steps and reduces the spike firing rate by more than 2×, significantly improving sparsity and energy efficiency without sacrificing expressivity.
@@ 3.3.3 Hardware Adaptation and Deployment Potential
However, the synchronous design of GPUs cannot fully exploit the event-driven and sparse-asynchronous advantages of spiking signals (schuman2022opportunities, ; frenkel2023bottom, ). GPUs operate under fixed high-frequency clock cycles, unlike biological neural systems that remain idle without spikes and trigger computation only upon spiking. To fully unlock the low-power potential of our scheme, deployment on specialized asynchronous hardware architectures—such as neuromorphic chips or spiking processors based on asynchronous circuit design—for matrix operations is required. These platforms can natively respond to sparse spike events without requiring clock synchronization (roy2019towards, ; Speck, ): circuits remain in a quiescent, low-power state in the absence of spikes and perform addition operations only when spikes occur.
@@ 3.4 Multi-stage Conversion Pipeline
The continual pre-training process comprises three stages which progressively extend the context window while adapting the model to new architectures. In the first stage, our models are trained on 100B tokens with a sequence length of 8k, aiming to transfer attention patterns toward local or low-rank variants and ensure loss convergence. Subsequently, the second and third stages extend the sequence length to 32k and 128k, respectively, each trained with 20B to 30B tokens. The entire conversion process consumes about 150B tokens in total. Compared to the $\sim$10T tokens typically required for training from scratch, the CPT approach requires only about 2% of the data, enabling efficient adaptation under resource and budget constraints. All three stages use the Matrix (zhang2024mapneo, ) dataset, with long-context data generated through a simple per-domain packing strategy. The RoPE base remains consistent with the base model at 1M.
SFT is conducted in three stages, each employing domain-specific data to progressively enhance the capabilities of the model in general knowledge, dialogues, and reasoning. The first stage focuses primarily on improving fundamental language understanding and domain knowledge, using the Infinity Instruct foundational (li2025infinity, ) dataset covering various foundational topics such as scientific knowledge, code interpretation, and mathematical problem solving. This stage is trained with 500k samples under an 8k sequence length. The second stage specializes in dialogue ability and instruction following, utilizing the Infinity Instruct chat (li2025infinity, ) dataset that includes multi-turn conversations, task-oriented dialogues, and knowledge-based question answering. The data volume and sequence length remain consistent with the first stage. The third stage targets reasoning tasks, using a high-quality reasoning dataset (Chinese-Data-Distill-From-R1, ; slam-distillation-from-r1, ) distilled via the DeepSeek-R1 (guo2025deepseek, ) method, containing examples with detailed chain-of-thought annotations for mathematical proofs, logical reasoning, case analyses, and other multi-step inference problems. To ensure cross-linguistic reasoning, the dataset maintains a 1:1 Chinese-English ratio. A total of 150k samples are used under a sequence length of 8k.

### v3

@@ 2.2 Integrated Model Architectures
The 7B model achieves purely linear complexity by interleaving linear attention and sliding window attention (SWA) layers with a fixed 4K window in a 1:1 ratio. The FFN modules adopt the same SwiGLU design as the base model. Spike coding is applied to convert the activations of all linear projection layers in the model, as matrix multiplications for parameter projections dominate the model’s FLOPs. In this architecture, the inter-layer hybrid design is optimized for extreme efficiency, owing to its simplicity in attention computation, model parallelism and sequence parallelism. SWA layers captures precise local patterns, while linear attention layers efficiently compresses long-range information. The model thus provides linear-time training complexity and constant memory usage during inference, regardless of sequence length, offering substantial efficiency gains for long-context processing. In practice, we implement linear attention using a Gated Linear Attention Module Yang et al. (2024c) (GLA), where the gating vector $\mathbf{g}_{t}$, derived from low-rank projection parameters, enhances expressivity and recurrent modeling capability:
@@
The 76B model integrates linear attention and SWA in a 1:1 intra-layer parallel hybrid configuration, while standard full-attention layers are interleaved at a 1:6 ratio across layers. This intra-layer hybrid design is optimized for stronger language modeling capabilities, as it enables more intricate attention computation and incorporates additional attention parameters and FLOPs. During parallel hybridization, outputs of both attention branches are normalized to ensure consistent scale and to prevent instability in the early stages of training. In addition, 128 learnable sink tokens (Xiao et al., 2023; Dong et al., 2024) are introduced to mitigate the attention-sink phenomenon in softmax attention and to enhance the flexibility of SWA for local modeling. Specifically, 128 trainable embeddings are prepended to the input; these tokens are attended by all other tokens and also attend to each other without causal masking. This functionality is implemented by customizing FlashAttention (Dao, 2024) kernels. The model also employs Gated Linear Attention Modules, but in this case, the gating vector is tied directly to the key vector (Chou et al., 2024; Qin et al., 2024), i.e., $\mathbf{g}_{t}=\mathbf{1}-\mathbf{k}_{t}$, eliminating the need for extra gating parameters. The FFN modules adopt a sparse Mixture-of-Experts (MoE) design: each MoE layer consists of 16 routed experts (top-1 activated) and one shared expert, so that only about 15% of parameters (12B) are activated per token. To stabilize training and control parameter growth, 7 dense FFN layers are preserved out of 28 total layers (specifically at layers 1, 2, 3, 5, 7, 9, and 11). Similarly, spike coding is applied to the activations of all linear projection layers throughout the model.
@@
Our architectural choices are closely aligned with principles observed in biological brains. i) Linear attention modules exhibit modeling properties analogous to human memory, relying on compressed and continuously updated states (Gu, 2025). At each step, they retrieve information only from the current memory, showing Markov-like behavior. From a biological perspective, their stateful temporal recurrence can be viewed as a simplified abstraction of dendritic dynamics with multi-branch morphology. ii) The Mixture-of-Experts (MoE) component embodies the principle of modular sparse activation and functional specialization, reminiscent of the distributed and specialized processing found in neural circuits (O’Doherty et al., 2021). iii) Our spike coding scheme draws inspiration from event-driven and adaptively sparse neuronal activation in biological systems (Maass, 1997). By combining network-level sparsity (MoE) with neuron-level spiking sparsity, our approach enables on-demand allocation of computation and provides a robust two-scale efficiency strategy. Collectively, these results suggest a promising pathway for designing large-model architectures that are both efficient and biologically plausible (Li et al., 2024).
@@ 3.1 Attention Map Correspondence
On this basis, SWA can be interpreted as a sparsified version of this map with a strong recency bias (Beltagy et al., 2020; Child et al., 2019; Zaheer et al., 2020):
@@
Moreover, linear attention can be regarded as a low-rank approximation of the attention map, where the maximum rank of $\mathbf{A}$ is $d$ (the dimensionality of the query/key vectors) (Chen et al., 2021; Wang et al., 2020):
@@
Leveraging this attention-map correspondence, we can initialize the QKV projection parameters of SpikingBrain directly from a pre-trained Transformer checkpoint (Kasai et al., 2021), and this applies to full attention, linear attention, and SWA. By reusing the learned QK similarity and performing lightweight training on a small amount of data, the attention can be adapted to these local or low-rank cases. Since SWA maintains local modeling precision and linear attention retains global interaction, combining them in a hybrid paradigm provides a closer approximation to the original attention map (Dass et al., 2023; Chen et al., 2021; Arora et al., 2024). This enables a smoother transition during conversion and accelerates loss convergence.
@@
  • Apply non-negative activation to the QK vectors in linear attention (Kasai et al., 2021; Mercat et al., 2024), such as ReLU or Sigmoid. Because softmax attention maps are inherently non-negative, the converted attention must preserve this property to enable smooth transfer. Normalization, as in softmax attention, is applied after the linear attention outputs.
@@
  • Fully train the model during conversion to ensure performance, using learning rate warm-up and either cross-architecture distillation (Wang et al., 2024; Zhang et al., 2024b) or full-parameter training (Zhang et al., 2024c). For simplicity and to fit the memory capacity of MetaX GPUs, we adopt full-parameter training without freezing the backbone.
@@ 3.2 MoE Upcycling
To efficiently expand a dense model into an MoE architecture, we adopt the upcycling technique (Komatsuzaki et al., 2022), which increases model capacity while reusing the knowledge encoded in the original parameters. At initialization, the FFN in the base dense model is replicated into $N$ experts, and a randomized router is introduced. The router selects the top-$k$ experts for each token with probability $\mathbf{p}$ and outputs their weighted sum, ensuring functional equivalence to the original dense FFN at the start:
@@
Directly replicating and activating multiple experts, however, amplifies the output scale. To maintain consistency between MoE and dense FFN outputs, we rescale the expert weights (He et al., 2024a). The relationship between MoE activation $a_{\text{MoE}}$ and dense activation $a_{\text{dense}}$ are derived as below, and the scaling factor $\alpha$ can be defined accordingly:
$a_{\text{MoE}}$ $=S\times a_{\text{dense}}+\frac{k}{N}\times a_{\text{dense}}=\left(S+\frac{k}{N}\right)a_{\text{dense}},\qquad\alpha=\sqrt[3]{\frac{1}{S+\frac{k}{N}}}.$ (18)
We apply this factor to both shared and routed experts at initialization, so that $E_{i}(\mathbf{x}):=\alpha\times\mathrm{FFN}$ holds for all $1\leq i\leq N$ at initialization.
@@ 3.3.1 Step 1: Adaptive Threshold Spiking
$\mathbf{v}_{t+1}=\mathbf{v}_{t}-V_{\text{th}}\cdot\mathbf{s}_{t}+\mathbf{x}_{t+1},\qquad\mathbf{s}_{t}=1\ \text{ if }\ \mathbf{v}_{t}\geq V_{\text{th}}(\mathbf{x})\ \text{ else }\ 0.$ (20)
@@
This single-step integer formulation enables computationally efficient and stable optimization on GPUs, while adaptive threshold maintains appropriate firing rates. As a result, our spiking scheme preserves the essential characteristics of biological neurons, yet is sufficiently simplified to be practical for large-scale models.
For gradient computation in the optimization phase, since the spiking process is an approximate representation of the original membrane potential, the gradient of the spiking process can be skipped (i.e., using straight-through estimation), which maintains the gradient computation process consistent with conventional training.
@@
  • Threshold dynamics: $V_{\text{th}}(\mathbf{x})$ is determined by the mean absolute membrane potential. When the input potential is high, $V_{\text{th}}(\mathbf{x})$ increases, preventing excessive spiking and thus controlling sparsity to reduce redundant computation. When the input potential is low, $V_{\text{th}}(\mathbf{x})$ decreases, allowing neurons to emit a small number of spikes to retain key information and avoid accuracy loss from inactivity. Statistically, membrane potentials often follow a long-tailed Gaussian-like distribution with occasional outliers. In this setting, the mean absolute value approximates 0.8 times the standard deviation, providing a stable metric for regulating spike activity.
  • Effect of the hyperparameter $k$: $k$ directly scales the threshold and thereby determines the distribution of spike counts. A larger $k$ lowers $V_{\text{th}}(\mathbf{x})$, producing higher spike counts $\mathbf{s}_{\text{INT}}$ and broader firing ranges. This is suited for accuracy-critical scenarios where additional computation is acceptable. A smaller $k$ raises $V_{\text{th}}(\mathbf{x})$, reducing spike counts and producing sparser encodings, which is advantageous for low-power edge deployments. Tuning $k$ thus provides a flexible trade-off between accuracy and efficiency.
@@ 3.3.2 Step 2: Spike Coding
During inference, the integer spike count $\mathbf{s}_{\text{INT}}$ generated in Step 1 is expanded into a sparse spike sequence with values $\{-1,0,1\}$ along the time dimension to support event-driven computation, which formulated as:
@@
  • Binary Spike Coding $\{0,1\}$: This is the most basic event-driven spike coding method. Each time a unit spike is fired (with a value of $1$), it represents an activation of the neuron state, and the spike count is accumulated over continuous time steps. This coding scheme is intuitive and simple, with low computational overhead, making it suitable for scenarios with very low spike counts and effectively reducing system complexity. However, when representing large counts, this coding often requires many time steps to complete the representation. Additionally, the lack of directional optimization for spike firing results in a higher firing rate, further limiting system energy efficiency. Meanwhile, this encoding scheme only supports positive integers (unsigned $\mathbf{s}_{\text{INT}}$) which is not applicable to Adaptive Threshold Spiking.
  • Ternary Spike Coding $\{-1,0,1\}$: To improve neuronal expressivity and sparsity in event-driven computations, we introduce ternary spike coding. The core idea is to extend the traditional binary spike coding by adding inhibitory spikes (1-1), resulting in three firing states: $\{-1,0,1\}$. Compared to binary coding, which can only represent positive values, ternary coding offers bidirectional expression capabilities. Here, "11" represents excitatory spikes, "1-1" inhibitory spikes, and "0" the silent state, making it more aligned with the "excitation/inhibition" regulatory mechanism in biological neural systems. As shown in Figure 4 (b), ternary spike coding not only provides symmetric expression capabilities of $\pm 1$ but also reconstructs the mapping from membrane potential to spike counts through a symmetric quantization strategy. This strategy maps activations to a symmetric distribution $k,\dots,0,\dots,+k]$ rather than $0,1,2,\dots]$, so that low-amplitude counts (e.g., $\pm 1$) statistically occupy a larger probability mass, effectively absorbing high-frequency large counts from the tail of the original distribution. As a result, this scheme halves the number of time steps and reduces the spike firing rate by more than 2×, significantly improving sparsity and energy efficiency without sacrificing expressivity.
@@ 3.3.3 Hardware Adaptation and Deployment Potential
However, the synchronous design of GPUs cannot fully exploit the event-driven and sparse-asynchronous advantages of spiking signals (Schuman et al., 2022; Frenkel et al., 2023). GPUs operate under fixed high-frequency clock cycles, unlike biological neural systems that remain idle without spikes and trigger computation only upon activation. To take full advantage of our design and leverage the low-power potential, implementing matrix operations on specialized asynchornous hardware architectures (e.g., neuromorphic chips, or spiking processors) is required. These platforms natively respond to sparse spike events without requiring clock synchronization (Roy et al., 2019; Yao et al., 2024b): circuits remain in a quiescent, low-power state in the absence of spikes and perform addition operations only when spikes occur.
@@ 3.4 Multi-stage Conversion Pipeline
The continual pre-training process comprises three stages which progressively extend the context window while adapting the model to new architectures. In the first stage, our models are trained on 100B tokens with a sequence length of 8k, aiming to transfer attention patterns toward local or low-rank variants and ensure loss convergence. Subsequently, the second and third stages extend the sequence length to 32k and 128k, respectively, each trained with 20B to 30B tokens. The entire conversion process consumes about 150B tokens in total. Compared to the $\sim$10T tokens typically required for training from scratch, the CPT approach requires only about 2% of the data, enabling efficient adaptation under resource and budget constraints. All three stages use the Matrix (Zhang et al., 2024a) dataset, with long-context data generated through a simple per-domain packing strategy. The RoPE base remains consistent with the base model at 1M.
SFT is conducted in three stages, each employing domain-specific data to progressively enhance the capabilities of the model in general knowledge, dialogues, and reasoning. The first stage focuses primarily on improving fundamental language understanding and domain knowledge, using the Infinity Instruct foundational (Li et al., 2025) dataset covering various foundational topics such as scientific knowledge, code interpretation, and mathematical problem solving. This stage is trained with 500k samples under an 8k sequence length. The second stage specializes in dialogue ability and instruction following, utilizing the Infinity Instruct chat (Li et al., 2025) dataset that includes multi-turn conversations, task-oriented dialogues, and knowledge-based question answering. The data volume and sequence length remain consistent with the first stage. The third stage targets reasoning tasks, using a high-quality reasoning dataset (Liu et al., 2025; Madhusudhan et al., 2025) distilled via the DeepSeek-R1 (Guo et al., 2025) method, containing examples with detailed chain-of-thought annotations for mathematical proofs, logical reasoning, case analyses, and other multi-step inference problems. To ensure cross-linguistic reasoning, the dataset maintains a 1:1 Chinese-English ratio. A total of 150k samples are used under a sequence length of 8k.